### v1

@@ 7.1 Radio-Quiet AGNs
AGNs are highly dynamic systems where physical components undergo continuous changes, giving rise to extreme variability phenomena. The X-ray flux from AGN exhibits stochastic, large-amplitude variations over a broad range of timescales, from years down to just a few minutes, the latter corresponding to the light-crossing time of one gravitational radius for typical local radio-quiet AGN [e.g. 360, 361, 362]. This variability is most commonly attributed to perturbations in the mass accretion rate, which modulate shorter-timescale fluctuations [e.g. 363]. However, additional factors, such as changes in coronal properties and sporadic absorption events [e.g. 364, 365, 366], can further contribute to flux variations. These mechanisms may even produce extreme variability events, in which the observed flux changes by one or more orders of magnitude on a timescale of months to years [e.g. 367, 368]. In this context, time-resolved spectral studies and spectral time analysis techniques [e.g. 369] are essential to unravel the physical origins of these fluctuations and unravel the various physical contributions to the variability of AGN. The large effective area of eXTP SFA will make such investigations possible with unprecedented accuracy.
@@
Over the past two decades, an increasing number of AGN have been observed to exhibit large-amplitude variability that cannot be easily explained by intrinsic local stochastic perturbations in the accretion rate. In many cases, these events appear to be driven by extrinsic phenomena, such as obscuration by low-ionization clumps (potentially originating from the dusty torus or the outer BLR) crossing our line of sight to the X-ray source [e.g. 367, 368]. In other cases, known as “changing-look AGN” (CLAGN, e.g. 370), extreme increases or drops in optical and X-ray flux, on timescales of months to years, are accompanied by the appearance or disappearance of the broad optical/UV emission lines. It is not yet clear if such transitions are driven by global changes in accretion rate or locally-operating instabilities such as the H-ionization [371] or radiation pressure [372] instabilities. Nonetheless, such events provide observational constraints on AGN duty cycles, and allow us to explore accretion flow state changes analogous to those occurring in BHXRBs [373].
@@
Because changes in the ionizing X-ray flux can have a significant impact on the BLR, fully understanding the CLAGN phenomenon requires not only constraining the contribution from the standard accretion disk but also tracking variations in the warm and hot Comptonization coronae. As discussed earlier, combined eXTP PFA and SFA observations will be crucial in constraining the two contributions, thus deciphering the intrinsic physics behind changing-look events. In particular, the large effective area and the ability to perform simultaneous spectro-polarimetric fits will enable precise detection of variations in key physical parameters and coronal geometry. At the same time X-ray spectral-timing analysis will offer critical insights into the mechanisms that drive these dramatic events, potentially revealing transitions in the accretion state of the source [e.g., 374].
@@
Rapid variations of the neutral absorption column density () on timescales ranging from a few hours to days are commonly observed in the X-ray spectra of bright, local, radio-quiet AGNs [e.g. 379, 380, 381, 341]. The most extreme cases of such variability manifest themselves as eclipses, where an obscuring cloud partially or fully blocks the X-ray source. Tracking these obscuration events provides critical insights into the structure, distance, and kinematics of the intervening material, offering a unique opportunity to probe the circumnuclear environment of AGN. The large effective area of SFA will enable high time-resolution spectral analysis, allowing the detection of these variations on the shortest timescales, which is essential for constraining the properties of clouds on the smallest orbits around the SMBH ([e.g. 382].
@@
Particularly valuable are observations that capture a complete eclipse, i.e. both the ingress and egress phases, as these provide measurements of cloud location, density, size, and even morphology [380, 383]. However, detecting such events remains challenging due to the difficulty in distinguishing eclipses from the intrinsic, short-timescale stochastic variability of AGN. As a result, only a limited number of full eclipses have been reported so far [384, 380, 385, 381, 341, 386]. The superior sensitivity of eXTP SFA will significantly increase the rate of eclipse detections, even for weak obscuration events. Figure 17 presents simulations of a 100 ks eXTP SFA light curve for a source like NGC 4051 (with a keV flux of ), undergoing a weak eclipse event (, and at maximum obscuration) occurring between 30 and 70 ks. The top-left panel of Figure 17 illustrates the variation in transmittance at different energies for a power-law continuum with photon index . This weak eclipse does not produce a clearly distinguishable dip in either the count rate or the softness ratio time series (middle-left panels). However, following the method of [378], the correlation between the softness ratio and count rate is fit with an empirical polynomial relation (Figure 17, right panel), and the best-fit model is subtracted from the softness ratio data. The lower-left panel of Figure 17 demonstrates that the eclipse signal becomes apparent in the residuals of the softness ratio.
@@
The ability to detect the weakest eclipses strongly depends on the signal-to-noise ratio of the light curves. Thanks to its large effective area, eXTP SFA will significantly improve the precision of this kind of measurements, enabling the detection of subtle eclipsing events that would otherwise remain hidden in the noise. Consequently, we can accumulate statistics on clouds’ locations and densities so we can test models of cloud distributions. Example model classes include ensembles of BLR clouds driving X-ray obscuration [387], or radiatively-driven winds that originate in the inner accretion flow and contain numerous partial-covering, X-ray-obscuring clumps [367, 388, 389]. This will allow for a more complete census of obscuration in AGN, ultimately refining our understanding of their circumnuclear environment.
@@
Beyond neutral absorption, AGN spectra also show evidence for the presence of powerful outflows which span a wide range of velocities, ionization states, and distances from the central engine. Such outflows carry momentum and energy, thus they might represent an important feedback mechanism in radio-quiet AGN, playing a crucial role in regulating the growth of both the SMBH and the host galaxy [390]. X-ray observations, particularly with the large effective area of SFA, will provide a unique window into the physics of ionized gas outflows, probing the innermost regions where the most energetic processes occur. In this context, ultrafast outflows (UFOs) are particularly intriguing, characterized by extreme velocities, typically 10%-30% of the speed of light, and detected as highly blueshifted absorption features primarily in the Fe K band [391]. Their high velocities and ionization parameters suggest that UFOs originate very close to the SMBH. The combination of high velocity and substantial column density implies that UFOs can carry significant accretion-driven kinetic energy, potentially playing a critical role in the AGN feedback [392]. The diagnostic power of eXTP SFA will be valuable, particularly for faint AGN with typical keV X-ray fluxes around erg cm-2 s-1. Simulations using a baseline model of a typical radio-quiet AGN – including a power-law continuum, a reflected component from neutral material (pexrav), and absorption by fast outflowing ionized gas (modeled using a grid of XSTAR photoionized absorption models) – show that with a 100 ks eXTP SFA exposure, it is possible to place tight constraints on the column density and ionization parameters of the absorber with low statistical uncertainties. Tracking the time-dependent ionization response of the winds to the rapidly variable continuum will allow us to constrain the density, location, 3-D structure, and kinematics of the winds, and thus pinpoint the wind launch mechanism and the wind acceleration zone [393, 394].
@@
X-ray variability has become apparent in IMBHs and can be considered to be a defining characteristic of these targets. The timescale of variability shown by IMBH candidates spans from minutes to hours on a short timescale and months to years on a long timescale [430]. The two well-known prototypical IMBH candidates, NGC 4395 and POX 52, are seen to be rapidly variable in X-rays, with strong amplitude variability [439, 440], consistent with scaled-down type 1 Seyferts. The high time resolution of s and large effective area provided by eXTP SFA will allow accurate characterization of the short-term variability of these sources. For rapidly variable targets, such as NGC 4395, the SFA’s long uninterrupted exposures will enable users to construct the PSD (power spectral density) over a broad frequency range, which is crucial for accurately identifying key features, such as characteristic break frequency timescales, and study their correlation with fundamental parameters, such as the BH mass and the mass accretion rate. Since this requires high signal-to-noise observations, the combination of enhanced sensitivity and large effective area of eXTP SFA will provide an opportunity to robustly examine the PSD. While single long exposures can be used to study the short-term flux variations, eXTP multi-epoch observations will enable accurate studies of the long-term X-ray variability, to investigate the underlying physical mechanisms driving these changes, thus shedding light on the structure of the accretion flow.
@@ 7.2 Radio-Loud AGNs
The emission from radio-loud (RL) AGNs spans nearly the entire electromagnetic spectrum and is predominantly attributed to radiation from their relativistic jets. The broadband spectral energy distributions (SEDs) of blazars typically exhibit a double-peaked structure. The first (lower frequency) peak, with peak frequency observed at infrared to X-ray wavelengths, and the second (higher frequency) peak, from MeV to GeV. The first peak is attributed to synchrotron radiation from relativistic electrons, while the origin of the second peak remains debated. Proposed mechanisms include inverse Compton (IC) scattering by relativistic electrons [e.g., 441, 442, 443, 444, 445] or emission from hadronic processes [e.g., 446, 447, 448].
Blazars are classified into two categories based on their emission line properties: flat spectrum radio quasars showing strong emission lines and BL Lac objects showing only weak emission lines or a featureless continuum [449]. A different classification into three categories has been proposed, based on the frequency of the synchrotron peak (): low-synchrotron-peaked blazars (LSPs; Hz), intermediate-synchrotron-peaked blazars (ISPs; Hz), and high-synchrotron-peaked blazars (HSPs; Hz) [450], see [451, 452]. The origin of X-ray emission varies across blazar classes. Tight constraints are expected to come from the investigation of multi-band intensity correlations, especially the correlations between the radio and the X-ray band, and the X-ray band and the TeV band. The detection of -ray emission from the large-scale jets of radio galaxies (RGs) [e.g., 453, 454, 455] confirms that some substructures also serve as sites for high-energy particle acceleration. However, the mechanisms governing particle acceleration and radiation, along with the composition of the jets, remain subjects of ongoing debate. Utilizing observations from eXTP and other multiwavelength detectors can provide valuable insights into the fundamental physics of relativistic jets in AGNs.
@@
Flares represent prominent observational phenomena during active phases of blazars. They are characterized by considerable flux variations over a wide range of timescales, from months to years [e.g. 457, 458], down to hours [e.g. 459, 460], or even minutes [e.g. 461, 462, 463]. Flares represent an important feature of blazar emission across the entire electromagnetic spectrum, which can provide direct insights into the microphysical processes in jets, such as particle acceleration mechanisms and energy dissipation. The physical origin of rapid flares remains debated, though these events are presumably due to strong and rapid energy dissipation in a relatively compact region within jet [e.g. 464, 465]. eXTP, with its high time resolution of s, will allow the detection and optimal sampling of the most extreme flux variations occurring in blazars. At the same time, eXTP’s spectral capabilities will enable dynamic tracking of spectral variations during these extreme flaring events, providing insights into the time-dependent evolution of the electron distribution. Moreover, by correlating the X-ray and -ray spectral properties and studying their evolution it will be possible to differentiate between SSC and EC contributions to the high-energy emission. Finally, time-dependent spectral modeling across different energy bands will place constraints on the location and physical properties of the emission regions.
eXTP SFA capabilities will also provide the opportunity to estimate the Doppler factor (). The -factor is an important parameter for radio loud AGNs, as it determines the amount of intrinsic flux enhancement and the apparent decrease of variability timescales. However, it can only be estimated using indirect methods [466, 467, 468]. As shown in [469], assuming that the X-ray and -ray emission come from the same emitting region, combined measurements of the X-ray spectrum and the corresponding variability timescale, can be used to estimate .
@@
Magnetic field structure, jet formation, and jet composition: X-ray polarization measurements can provide a new perspective for studying the properties of relativistic jets, including radiation mechanisms, particle acceleration, macro/micro dynamics, magnetic field topology, and geometry of emission region. For instance, the Imaging X-ray Polarimetry Explorer (IXPE) observed a progressive decrease in the degree of polarization from X-rays to optical and radio wavelengths in Mrk 501 [470], which has been associated with a shock front as the origin of particle acceleration. This result highlights the importance of X-ray observations for our understanding of jet physics and structure. The magnetic-driven model posits that the rotational energy of the black hole accretion system is converted into Poynting flux, which subsequently accelerates the jet [471, 472, 473]. According to this model, a helical magnetic field plays a crucial role in the jet’s formation, collimation, and acceleration [473]. However, observational evidence supporting this theory remains limited. IXPE observations revealed polarization angle rotation in Mkn 421 by 360 degrees, hinting at the presence of a helical magnetic field in this source [474]. However, this evidence is still inconclusive, emphasizing the need for further investigations with enhanced sensitivity and high-quality data. Targeted studies of High-Synchrotron-Peaked (HSP) blazars using the eXTP-PFA will be crucial in providing deep insights into the magnetic field configuration of AGN jets, thereby advancing our understanding of the mechanisms that drive jet formation, collimation, and acceleration. In addition, thanks to its sensitivity, eXTP-SFA, together with other multi-wavelength facilities, will allow the SED to be constrained, in order to properly infer important physical parameters, such as the electron energy (), the magnetic field strength (), and the -ray emission location [475].
@@
The unprecedented high sensitivity measurements of X-ray spectra and polarization in RLAGNs with eXTP will be key to determine whether hadronic models should be invoked to explain the observed emission. In particular, in the case of the IC process, X-ray polarization is expected to be significantly lower than that of the seed photons [476, 477, 478, 479, 480]. Conversely, for hadronic interactions, X-ray polarization should be similar to or even higher than that observed in radio or optical wavelengths (e.g., in pure proton-synchrotron models) [481, 482, 483].
@@
Polarization simulation: eXTP-PFA capabilities will also enable unprecedented energy-dependent polarimetric studies. To demonstrate this, we consider the BL Lac object, 1ES 1959+650, with X-ray flux of about [484], and linear polarization degree of as constrained by IXPE [485]. Simulations of the expected as a function of energy for 1ES 1959+650 are shown in Figure 18 for different exposure times, demonstrating PFA’s unprecedented sensitivity to polarization. Here, MDP means the minimum detectable polarization.
@@
The X-ray emission of some substructures in large-scale jets is thought to originate from synchrotron radiation [e.g., 486, 468]. However, this emission is apparently not a simple extrapolation of the radio-to-optical synchrotron spectrum and requires an additional component for explanation. Such component could either arise from IC scattering [487] or synchrotron radiation from a distinct population of high-energy particles [e.g., 488, 489]. High sensitivity X-ray polarization observations will be a powerful tool to elucidate the X-ray radiation mechanism, particularly for the substructures with highly ordered magnetic fields, such as the western hotspot (WHS) in RG Pictor A [e.g., 486, 490]. The WHS in Pictor A is a faint emission region with X-ray flux of . According to eXTP-PFA simulations, the will be about 30% and 60% in the energy bands of keV and keV, respectively. Although the polarization measurement for the detailed structure in a global jet is important to investigate the magnetic field and dynamical properties in a non-uniform jet, it is a challenge for the eXTP polarization to detect such weak emission.
@@
Polarized flares as detected by eXTP-PFA: Rapid changes in polarization degree during a blazar flare may indicate localized energy release and particle acceleration, possibly triggered by the radiative magnetic reconnection or turbulence [491, 492, 493, 494, 495]. In particular, it is interesting to investigate the critical Synchrotron/SSC transition regions of ISPs. During flaring states, the synchrotron peak may shift to higher energy bands, leading to changes in X-ray emission polarization as flux varies, transitioning from unpolarized to significantly polarized states. Energy-dependent polarization analysis during these transitions may reveal the Synchrotron/SSC transition regions within the broadband SED of ISPs [496]. According to simulations shown in Figure 18, eXTP-PFA will be able to perform such polarization measurements at least for the brightest blazars (with X-ray flux of ). At the same time, a comparative analysis of polarization variations across different wavebands can shed light on the energy dissipation across different timescales.

### v2

@@ 7.1 Radio-Quiet AGNs
AGNs are highly dynamic systems where physical components undergo continuous changes, giving rise to extreme variability phenomena. The X-ray flux from AGN exhibits stochastic, large-amplitude variations over a broad range of timescales, from years down to just a few minutes, the latter corresponding to the light-crossing time of one gravitational radius for typical local radio-quiet AGN [e.g. 365, 366, 367]. This variability is most commonly attributed to perturbations in the mass accretion rate, which modulate shorter-timescale fluctuations [e.g. 368]. However, additional factors, such as changes in coronal properties and sporadic absorption events [e.g. 369, 370, 371], can further contribute to flux variations. These mechanisms may even produce extreme variability events, in which the observed flux changes by one or more orders of magnitude on a timescale of months to years [e.g. 372, 373]. In this context, time-resolved spectral studies and spectral time analysis techniques [e.g. 374] are essential to unravel the physical origins of these fluctuations and unravel the various physical contributions to the variability of AGN. The large effective area of eXTP SFA will make such investigations possible with unprecedented accuracy.
@@
Over the past two decades, an increasing number of AGN have been observed to exhibit large-amplitude variability that cannot be easily explained by intrinsic local stochastic perturbations in the accretion rate. In many cases, these events appear to be driven by extrinsic phenomena, such as obscuration by low-ionization clumps (potentially originating from the dusty torus or the outer BLR) crossing our line of sight to the X-ray source [e.g. 372, 373]. In other cases, known as “changing-look AGN” (CLAGN, e.g. 375), extreme increases or drops in optical and X-ray flux, on timescales of months to years, are accompanied by the appearance or disappearance of the broad optical/UV emission lines. It is not yet clear if such transitions are driven by global changes in accretion rate or locally-operating instabilities such as the H-ionization [376] or radiation pressure [377] instabilities. Nonetheless, such events provide observational constraints on AGN duty cycles, and allow us to explore accretion flow state changes analogous to those occurring in BHXRBs [378].
@@
Because changes in the ionizing X-ray flux can have a significant impact on the BLR, fully understanding the CLAGN phenomenon requires not only constraining the contribution from the standard accretion disk but also tracking variations in the warm and hot Comptonization coronae. As discussed earlier, combined eXTP-PFA and SFA observations will be crucial in constraining the two contributions, thus deciphering the intrinsic physics behind changing-look events. In particular, the large effective area and the ability to perform simultaneous spectro-polarimetric fits will enable precise detection of variations in key physical parameters and coronal geometry. At the same time X-ray spectral-timing analysis will offer critical insights into the mechanisms that drive these dramatic events, potentially revealing transitions in the accretion state of the source [e.g., 379].
@@
Rapid variations of the neutral absorption column density () on timescales ranging from a few hours to days are commonly observed in the X-ray spectra of bright, local, radio-quiet AGNs [e.g. 384, 385, 386, 346]. The most extreme cases of such variability manifest themselves as eclipses, where an obscuring cloud partially or fully blocks the X-ray source. Tracking these obscuration events provides critical insights into the structure, distance, and kinematics of the intervening material, offering a unique opportunity to probe the circumnuclear environment of AGN. The large effective area of SFA will enable high time-resolution spectral analysis, allowing the detection of these variations on the shortest timescales, which is essential for constraining the properties of clouds on the smallest orbits around the SMBH ([e.g. 387].
@@
Particularly valuable are observations that capture a complete eclipse, i.e. both the ingress and egress phases, as these provide measurements of cloud location, density, size, and even morphology [385, 388]. However, detecting such events remains challenging due to the difficulty in distinguishing eclipses from the intrinsic, short-timescale stochastic variability of AGN. As a result, only a limited number of full eclipses have been reported so far [389, 385, 390, 386, 346, 391]. The superior effective area of eXTP-SFA will significantly increase the rate of eclipse detections, even for weak obscuration events. Figure 17 presents simulations of a 100 ks eXTP-SFA light curve for a source like NGC 4051 (with a keV flux of ), undergoing a weak eclipse event (, and at maximum obscuration) occurring between 30 and 70 ks. The top-left panel of Figure 17 illustrates the variation in transmittance at different energies for a power-law continuum with photon index . This weak eclipse does not produce a clearly distinguishable dip in either the count rate or the softness ratio time series (middle-left panels). However, following the method of [383], the correlation between the softness ratio and count rate is fit with an empirical polynomial relation (Figure 17, right panel), and the best-fit model is subtracted from the softness ratio data. The lower-left panel of Figure 17 demonstrates that the eclipse signal becomes apparent in the residuals of the softness ratio.
@@
The ability to detect the weakest eclipses strongly depends on the signal-to-noise ratio of the light curves. Thanks to its large effective area, eXTP-SFA will significantly improve the precision of this kind of measurements, enabling the detection of subtle eclipsing events that would otherwise remain hidden in the noise. Consequently, we can accumulate statistics on clouds’ locations and densities so we can test models of cloud distributions. Example model classes include ensembles of BLR clouds driving X-ray obscuration [392], or radiatively-driven winds that originate in the inner accretion flow and contain numerous partial-covering, X-ray-obscuring clumps [372, 393, 394]. This will allow for a more complete census of obscuration in AGN, ultimately refining our understanding of their circumnuclear environment.
@@
Beyond neutral absorption, AGN spectra also show evidence for the presence of powerful outflows which span a wide range of velocities, ionization states, and distances from the central engine. Such outflows carry momentum and energy, thus they might represent an important feedback mechanism in radio-quiet AGN, playing a crucial role in regulating the growth of both the SMBH and the host galaxy [395]. X-ray observations, particularly with the large effective area of SFA, will provide a unique window into the physics of ionized gas outflows, probing the innermost regions where the most energetic processes occur. In this context, ultrafast outflows (UFOs) are particularly intriguing, characterized by extreme velocities, typically 10%-30% of the speed of light, and detected as highly blueshifted absorption features primarily in the Fe K band [396]. Their high velocities and ionization parameters suggest that UFOs originate very close to the SMBH. The combination of high velocity and substantial column density implies that UFOs can carry significant accretion-driven kinetic energy, potentially playing a critical role in the AGN feedback [397]. The UFOs from AGN could produce gamma-ray emission [398], which is also a proof of carrying significant kinetic energy. The diagnostic power of eXTP-SFA will be valuable, particularly for faint AGN with typical keV X-ray fluxes around erg cm-2 s-1. Simulations using a baseline model of a typical radio-quiet AGN – including a power-law continuum, a reflected component from neutral material (pexrav), and absorption by fast outflowing ionized gas (modeled using a grid of XSTAR photoionized absorption models) – show that with a 100 ks eXTP-SFA exposure, it is possible to place tight constraints on the column density and ionization parameters of the absorber with low statistical uncertainties. Tracking the time-dependent ionization response of the winds to the rapidly variable continuum will allow us to constrain the density, location, 3-D structure, and kinematics of the winds, and thus pinpoint the wind launch mechanism and the wind acceleration zone [399, 400].
@@
X-ray variability has become apparent in IMBHs and can be considered to be a defining characteristic of these targets. The timescale of variability shown by IMBH candidates spans from minutes to hours on a short timescale and months to years on a long timescale [436]. The two well-known prototypical IMBH candidates, NGC 4395 and POX 52, are seen to be rapidly variable in X-rays, with strong amplitude variability [445, 446], consistent with scaled-down type 1 Seyferts. The high time resolution of s and large effective area provided by eXTP SFA will allow accurate characterization of the short-term variability of these sources. For rapidly variable targets, such as NGC 4395, the SFA’s long uninterrupted exposures will enable users to construct the PSD (power spectral density) over a broad frequency range, which is crucial for accurately identifying key features, such as characteristic break frequency timescales, and study their correlation with fundamental parameters, such as the BH mass and the mass accretion rate. Since this requires high signal-to-noise observations, the large effective area of eXTP-SFA will provide an opportunity to robustly examine the PSD. While single long exposures can be used to study the short-term flux variations, eXTP multi-epoch observations will enable accurate studies of the long-term X-ray variability, to investigate the underlying physical mechanisms driving these changes, thus shedding light on the structure of the accretion flow.
@@ 7.2 Radio-Loud AGNs
The emission from radio-loud (RL) AGNs spans nearly the entire electromagnetic spectrum and is predominantly attributed to radiation from their relativistic jets. The broadband spectral energy distributions (SEDs) of blazars typically exhibit a double-peaked structure. The first (lower frequency) peak, with peak frequency observed at infrared to X-ray wavelengths, and the second (higher frequency) peak, from MeV to GeV. The first peak is attributed to synchrotron radiation from relativistic electrons, while the origin of the second peak remains debated. Proposed mechanisms include inverse Compton (IC) scattering by relativistic electrons [e.g., 447, 448, 449, 450, 451] or emission from hadronic processes [e.g., 452, 453, 454].
Blazars are classified into two categories based on their emission line properties: flat spectrum radio quasars showing strong emission lines and BL Lac objects showing only weak emission lines or a featureless continuum [455]. A different classification into three categories has been proposed, based on the frequency of the synchrotron peak (): low-synchrotron-peaked blazars (LSPs; Hz), intermediate-synchrotron-peaked blazars (ISPs; Hz), and high-synchrotron-peaked blazars (HSPs; Hz) [456], see [457, 458]. The origin of X-ray emission varies across blazar classes. Tight constraints are expected to come from the investigation of multi-band intensity correlations, especially the correlations between the radio and the X-ray band, and the X-ray band and the TeV band. The detection of -ray emission from the large-scale jets of radio galaxies (RGs) [e.g., 459, 460, 461] confirms that some substructures also serve as sites for high-energy particle acceleration. However, the mechanisms governing particle acceleration and radiation, along with the composition of the jets, remain subjects of ongoing debate. Utilizing observations from eXTP and other multiwavelength detectors can provide valuable insights into the fundamental physics of relativistic jets in AGNs.
@@
Flares represent prominent observational phenomena during active phases of blazars. They are characterized by considerable flux variations over a wide range of timescales, from months to years [e.g. 463, 464], down to hours [e.g. 465, 466], or even minutes [e.g. 467, 468, 469]. Flares represent an important feature of blazar emission across the entire electromagnetic spectrum, which can provide direct insights into the microphysical processes in jets, such as particle acceleration mechanisms and energy dissipation. The physical origin of rapid flares remains debated, though these events are presumably due to strong and rapid energy dissipation in a relatively compact region within jet [e.g. 470, 471]. eXTP, with its high time resolution of s, will allow the detection and optimal sampling of the most extreme flux variations occurring in blazars. At the same time, eXTP’s spectral capabilities will enable dynamic tracking of spectral variations during these extreme flaring events, providing insights into the time-dependent evolution of the electron distribution. Moreover, by correlating the X-ray and -ray spectral properties and studying their evolution it will be possible to differentiate between SSC and EC contributions to the high-energy emission. Finally, time-dependent spectral modeling across different energy bands will place constraints on the location and physical properties of the emission regions.
eXTP-SFA will also provide the opportunity to estimate the Doppler factor (). The -factor is an important parameter for radio loud AGNs, as it determines the amount of intrinsic flux enhancement and the apparent decrease of variability timescales. However, it can only be estimated using indirect methods [472, 473, 474]. As shown in [475], assuming that the X-ray and -ray emission come from the same emitting region, combined measurements of the X-ray spectrum and the corresponding variability timescale, can be used to estimate .
@@
Magnetic field structure, jet formation, and jet composition: X-ray polarization measurements can provide a new perspective for studying the properties of relativistic jets, including radiation mechanisms, particle acceleration, macro/micro dynamics, magnetic field topology, and geometry of emission region. For instance, the Imaging X-ray Polarimetry Explorer (IXPE) observed a progressive decrease in the degree of polarization from X-rays to optical and radio wavelengths in Mrk 501 [476], which has been associated with a shock front as the origin of particle acceleration. This result highlights the importance of X-ray observations for our understanding of jet physics and structure. The magnetic-driven model posits that the rotational energy of the black hole accretion system is converted into Poynting flux, which subsequently accelerates the jet [477, 478, 479]. According to this model, a helical magnetic field plays a crucial role in the jet’s formation, collimation, and acceleration [479]. However, observational evidence supporting this theory remains limited. IXPE observations revealed polarization angle rotation in Mkn 421 by 360 degrees, hinting at the presence of a helical magnetic field in this source [480]. However, this evidence is still inconclusive, emphasizing the need for further investigations with enhanced sensitivity and high-quality data. Targeted studies of High-Synchrotron-Peaked (HSP) blazars using the eXTP-PFA will be crucial in providing deep insights into the magnetic field configuration of AGN jets, thereby advancing our understanding of the mechanisms that drive jet formation, collimation, and acceleration. In addition, thanks to its larger effective area, eXTP-SFA, together with other multi-wavelength facilities, will allow the SED to be constrained, in order to properly infer important physical parameters, such as the electron energy (), the magnetic field strength (), and the -ray emission location [481].
@@
With a large effective area, eXTP will provide high-quality X-ray spectral and polarization measurements for RLAGNs, which is key to determining whether hadronic models should be invoked to explain the observed emission. In particular, in the case of the IC process, X-ray PD is expected to be significantly lower than that of the seed photons [482, 483, 484, 485, 486]. Conversely, for hadronic interactions, X-ray PD should be similar to or even higher than that observed in radio or optical wavelengths (e.g., in pure proton-synchrotron models) [487, 488, 489].
@@
eXTP-PFA will enable unprecedented energy-dependent polarimetric studies. To demonstrate this, we consider the BL Lac object, 1ES 1959+650, with X-ray flux of about [490], and linear polarization degree of as constrained by IXPE [491]. Simulations of the expected as a function of energy for 1ES 1959+650 are shown in Figure 18 for different exposure times, demonstrating PFA’s unprecedented sensitivity to polarization. Here, MDP means the minimum detectable polarization.
@@
The X-ray emission of some substructures in large-scale jets is thought to originate from synchrotron radiation [e.g., 492, 474]. However, this emission is apparently not a simple extrapolation of the radio-to-optical synchrotron spectrum and requires an additional component for explanation. Such component could either arise from ICS [493] or synchrotron radiation from a distinct population of high-energy particles [e.g., 494, 495]. High sensitivity X-ray polarization observations will be a powerful tool to elucidate the X-ray radiation mechanism, particularly for the substructures with highly ordered magnetic fields, such as the western hotspot (WHS) in RG Pictor A [e.g., 492, 496]. The WHS in Pictor A is a faint emission region with X-ray flux of . According to eXTP-PFA simulations, the will be about 30% and 60% in the energy bands of keV and keV, respectively. Although the polarization measurement for the detailed structure in a global jet is important to investigate the magnetic field and dynamical properties in a non-uniform jet, it is a challenge for the eXTP polarization to detect such weak emission.
@@
Rapid changes in polarization degree during a blazar flare may indicate localized energy release and particle acceleration, possibly triggered by the radiative magnetic reconnection or turbulence [497, 498, 499, 500, 501]. In particular, it is interesting to investigate the critical Synchrotron/SSC transition regions of ISPs. During flaring states, the synchrotron peak may shift to higher energy bands, leading to changes in X-ray emission polarization as flux varies, transitioning from unpolarized to significantly polarized states. Energy-dependent polarization analysis during these transitions may reveal the Synchrotron/SSC transition regions within the broadband SED of ISPs [502]. According to simulations shown in Figure 18, eXTP-PFA will be able to perform such polarization measurements at least for the brightest blazars (with X-ray flux of ). At the same time, a comparative analysis of polarization variations across different wavebands can shed light on the energy dissipation across different timescales.